OpenStudy (anonymous):

x^2+4x-50 divided by x+7=

OpenStudy (saifoo.khan):

Quotient = x-3 Remainder = -29

OpenStudy (saifoo.khan):

If you are doing this online, so it will be written as: \[x-3 +\frac{-29}{x+7}\]